{3x+2y=19}
{x+2y=9}

Answer by london maths tutor(243)

{3x+2y=19}... equation 1
{x+2y=9} ... equation 2
equation 1 minus equation 2
3x + 2y - (x + 2y)= 19 - 9
3x - x + 2y - 2y = 10
2x = 10
x = 10/2
x = 5
therefore, (5)+ 2y = 9
2y = 9 - 5
2y = 4
y = 4/2
y = 2
Answer: x = 5 and y = 2